# Linear Complexity Gibbs Sampling for Generalized Labeled Multi-Bernoulli Filtering

### v1

@@ III-B Tempered Gibbs Sampling
Similar to RGS, TGS also generates the next iterate $\gamma^{\prime}_{\text{+}}$ by randomly selecting a coordinate to update. However, TGS provides an additional mechanism to improve mixing and sample diversity [38]. Specifically, a coordinate $i\in\{1\text{:}P\}$ is chosen according to the distribution
@@
[…]
$\gamma^{\prime}_{\text{+}}(\ell_{i})\sim\phi_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}})).$ (19)
The proposal $\phi_{i}$ controls sample diversity, and determines coordinate selection that can influence mixing [44, 45, 42]. Note that TGS reduces to RGS in the special case $\phi_{i}$ = $\pi_{i}$.
In GLMB truncation, we are not interested in importance weights since the goal is to generate distinct samples with significant GLMB weights (which are different from the importance weights). While TGS can circumvent the drawbacks of RGS by using fast mixing proposals that yield diverse samples, these may not be significant (in GLMB weights) because the tempered stationary distribution could be very different from $\pi$. One way to generate diverse and significant samples is to use proposals that approximate the conditional $\pi_{i}$, but are more diffuse, which can be achieved with a mixture consisting of the conditional and its tempered version, i.e.,
@@
Input : $\gamma_{\text{+}}$, $\alpha$, $\beta$, $[\eta_{i}(j)]_{i\text{=}1}^{P}$
@@
  • Computing the proposals $\phi_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))$ for all $i\in\{1\mathnormal{:}P\}$ incurs $\mathcal{O}(PM)$ complexity due to $\beta$-th power operations and normalizations;
@@ III-C Linear Complexity TGS
where $\nu_{i}^{(\beta)}(\gamma_{\text{+}})=\left\langle\widetilde{\pi}_{i}^{\beta}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}})),1\right\rangle$ is the normalizing constant for $\widetilde{\pi}_{i}^{\beta}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))$ (hence $\nu_{i}^{(1)}(\gamma_{\text{+}})$ is the normalizing constant for $\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))$). Further, the unnormalized conditional $\widetilde{\pi}_{i}(\cdot|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))$ at the next GS iteration differs from $\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))$ at no more than two points on its domain {-$\texttt{-}1\mathnormal{:}M$}. Fig. 2 provides an illustration.
Figure 2: Difference between unnormalized conditionals in successive TGS iterations. If $\gamma^{\prime}_{\text{+}}$ differs from $\gamma_{\text{+}}$ only at the $n$-th coordinate, i.e., $\gamma_{\text{+}}(\ell_{n})=j>0$ and $\gamma^{\prime}_{\text{+}}(\ell_{n})=j^{\prime}>0$, with $j\neq j^{\prime}$, then for $i\in\{\bar{n}\}$, $\widetilde{\pi}_{i}(\cdot|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))$ is the same as $\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))$ on {-$\{\texttt{-}1\mathnormal{:}M\}$ except at $j$ (gray) and $j^{\prime}$ (pink).
@@
Suppose that $\gamma^{\prime}_{\text{+}}$ and $\gamma_{\text{+}}$ differ only at the $n$-th coordinate, i.e., $\gamma^{\prime}_{\text{+}}(\ell_{\bar{n}})=\gamma_{\text{+}}(\ell_{\bar{n}})$ and $\gamma^{\prime}_{\text{+}}(\ell_{n})\neq\gamma_{\text{+}}(\ell_{n})$. Then, for $\beta>0$,
@@
Proof:
(i) For $j<1$, it follows from Proposition 1 that $\widetilde{\pi}_{i}(j|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))=\eta_{i}(j)=\widetilde{\pi}_{i}(j|\gamma_{\text{+}}(\ell_{\bar{i}}))$.
(ii) For $j>0$, note that $n\in\{\bar{i}\}$ implies $\gamma_{\text{+}}(\ell_{n})\in\{\gamma_{\text{+}}(\ell_{\bar{i}})\}$, and $\gamma^{\prime}_{\text{+}}(\ell_{n})\in\{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})\}$, i.e.,
$1_{\gamma_{\text{+}}(\ell_{\bar{i}})}(\gamma_{\text{+}}(\ell_{n}))$ $=1,$ (22)
$1_{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})}(\gamma^{\prime}_{\text{+}}(\ell_{n}))$ $=1;$ (23)
and hence $\gamma_{\text{+}}(\ell_{n})\notin\{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})\}$, $\gamma^{\prime}_{\texttt{+}}(\ell_{n})\notin\{\gamma_{\texttt{+}}(\ell_{\bar{i}})\}$ i.e.,
Now, apply Proposition 1 to each of the following cases.
(a) $j=\gamma_{\text{+}}(\ell_{n})$:
$\widetilde{\pi}_{i}(j|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))=\eta_{i}(j)(1-1_{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})}(j))=\eta_{i}(j),$ using (26);
(b) $j=\gamma^{\prime}_{\text{+}}(\ell_{n})$:
$\widetilde{\pi}_{i}(j|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))=\eta_{i}(j)(1-1_{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})}(j))=0,$ using (23);
$\widetilde{\pi}_{i}(j|\gamma_{\text{+}}(\ell_{\bar{i}}))=\eta_{i}(j)(1-1_{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})}(j))=\eta_{i}(j),$ using (27).
(c) $j\neq\gamma^{\prime}_{\text{+}}(\ell_{n})$ and $j\neq\gamma_{\text{+}}(\ell_{n})$: it follows from (24), (25) that $j\in\{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})\}$ iff $j\in\{\gamma_{\text{+}}(\ell_{\bar{i}})\}$, i.e., $1_{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})}(j)=1_{\gamma_{\text{+}}(\ell_{\bar{i}})}(j)$. Hence $\widetilde{\pi}_{i}(j|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))=\eta_{i}(j)(1-1_{\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}})}(j))=\widetilde{\pi}_{i}(j|\gamma_{\text{+}}(\ell_{\bar{i}}))$.
Raising the unnormalized conditionals to the power of $\beta>0,$ and taking their difference, i.e., $\widetilde{\pi}_{i}^{\beta}(j|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))-\widetilde{\pi}_{i}^{\beta}(j|\gamma_{\text{+}}(\ell_{\bar{i}}))$, gives the desired result. ∎
@@
Given successive TGS iterates $\gamma_{\text{+}}$ and $\gamma^{\prime}_{\text{+}}$ that differ at the $n$-th coordinate, Proposition 4 (with $\beta=1$) states that their $n$-th conditionals are the same, i.e., $\pi_{n}(\cdot|\gamma_{\text{+}}(\ell_{\bar{n}}))=\pi_{n}(\cdot|\gamma^{\prime}_{\text{+}}(\ell_{\bar{n}}))$. Moreover, for $i\neq n$ the unnormalized $i$-th conditional $\widetilde{\pi}_{i}(\cdot|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))$ is the same as $\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))$, except at:
[…]
@@
For successive TGS iterates $\gamma_{\text{+}}$ and $\gamma^{\prime}_{\text{+}}$ differing at the $n$-th coordinate,
@@
$+$ $\eta_{i}^{\beta}(\gamma_{\text{+}}(\ell_{n}))1_{\{1:M\}}(\gamma_{\text{+}}(\ell_{n}))$
-- $\eta_{i}^{\beta}(\gamma^{\prime}_{\text{+}}(\ell_{n}))1_{\{1:M\}}(\gamma^{\prime}_{\text{+}}(\ell_{n})).$
@@
The above result means propagating each tempered conditional and its normalizing constant to the next iterate of the Markov chain can be performed with a constant time complexity. Consequently, $\mathcal{O}(T(P+M))$ complexity TGS can be developed for selecting significant GLMB components. The pseudocode in Algorithm 2, herein referred to as $TGS^{+}$, outlines the steps for generating iterates $\gamma_{\text{+}}^{(1)},...,\gamma_{\text{+}}^{(T)}$, from initial chain state $\gamma_{\text{+}}^{(0)}$, and initial coordinate distribution $\rho(\cdot|\gamma_{\text{+}}^{(0)})$. Note that:
@@
  • Sampling $n$ from the categorical distribution $\rho(\cdot|\gamma_{\text{+}}^{(t\texttt{-}1)})$, defined on $P$ categories, incurs $\mathcal{O}(P)$ complexity (line 2);
@@
  • Computing the proposal $\phi_{n}(\cdot|\gamma_{\text{+}}^{(t\texttt{-}1)}(\ell_{\bar{n}}))$ incurs $\mathcal{O}(M)$ complexity (line 2) since $\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}^{(t\texttt{-}1)}(\ell_{\bar{i}}))$, $\nu_{i}^{(1)}(\gamma_{\text{+}}^{(t\texttt{-}1)})$, and $\nu_{i}^{(\beta)}(\gamma_{\text{+}}^{(t\texttt{-}1)})$ were generated as by-products of computing $\rho(\cdot|\gamma_{\text{+}}^{(t\texttt{-}1)})$ in the previous iteration;
@@
  • Sampling $\gamma_{\text{+}}^{(t)}(\ell_{n})$ from $\phi_{n}(\cdot|\gamma_{\text{+}}^{(t\texttt{-}1)}(\ell_{\bar{n}}))$ incurs $\mathcal{O}(M)$ complexity (line 2);
@@
  • Computing $\widetilde{\rho}(\cdot|\gamma_{\text{+}}^{(t)})$ for the $(t+1)$-th iteration is only needed when $\gamma_{\text{+}}^{(t)}(\ell_{n})\neq\gamma_{\text{+}}^{(t\texttt{-}1)}(\ell_{n})$, and requires $\mathcal{O}(P)$ complexity (lines 2-2), because for each $i\in\{1\mathnormal{:}P\}$, evaluating (18) from the available $\pi_{i}(\gamma_{\text{+}}^{(t)}(\ell_{i})|\gamma_{\text{+}}^{(t)}(\ell_{\bar{i}}))$ and $\phi_{i}(\gamma_{\text{+}}^{(t)}(\ell_{i})|\gamma_{\text{+}}^{(t)}(\ell_{\bar{i}}))$ only requires a constant time complexity (Corollary 6);

### v2

@@ III-B Tempered Gibbs Sampling
Similar to RGS, TGS also generates the next iterate $\gamma^{\prime}_{\text{+}}$ by randomly selecting a coordinate to update. However, TGS provides an additional mechanism to improve mixing and sample diversity[38]. Specifically, a coordinate $i\in\{1\mathnormal{:}P\}$ is chosen according to the distribution
@@
where $\phi_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))$ is a bounded proposal, defined on {-$\{\texttt{-}1\text{:}M\}$, with the same support as $\pi_{i}(\cdot|\gamma(\ell_{\bar{i}}))$. Further, given the selection of the $i$-th coordinate, its state is updated by sampling from the proposal, i.e.,
$\gamma^{\prime}_{\text{+}}(\ell_{i})\sim\phi_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}})),$ (20)
(note that TGS reduces to RGS in the special case $\phi_{i}=\pi_{i}$). The proposal $\phi_{i}$ controls sample diversity, and determines coordinate selection that can influence mixing[48, 49, 46]. The transition kernel of TGS is given by $\phi(\gamma^{\prime}_{\text{+}}|\gamma_{\text{+}})\propto\rho(i)\phi_{i}(\gamma^{\prime}_{\text{+}}(\ell_{i})|\gamma_{\text{+}}(\ell_{\bar{i}}))$ when $\gamma_{\text{+}}$ and $\gamma^{\prime}_{\text{+}}$ differ at most in the $i$-th coordinate and 00 otherwise.
In GLMB truncation, we are not interested in the importance weights since the goal is to generate distinct samples with significant GLMB weights (which are different from the importance weights). While TGS can circumvent the drawbacks of RGS by using fast mixing proposals that yield diverse samples, these may not be significant (in GLMB weights) because the tempered stationary distribution could be very different from $\pi$. One way to generate diverse and significant samples is to use proposals that approximate the conditional $\pi_{i}$, but are more diffuse, which can be achieved with a mixture consisting of the conditional and its tempered version, i.e.,
@@
Input : $\gamma_{\text{+}}$, $\beta$, $[\eta_{i}(j)]^{P}_{i\texttt{=}1}$
@@
  • Computing the proposals $\phi_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))$ for all $i\in\{1\mathnormal{:}P\}$ incurs $\mathcal{O}(PM)$ complexity due to $\beta$-th power operations and normalizations; and
@@ III-C Linear Complexity TGS
where $\nu_{i}^{(\beta)}(\gamma_{\text{+}})=\left\langle\widetilde{\pi}_{i}^{\beta}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}})),1\right\rangle$ is the normalizing constant for $\widetilde{\pi}_{i}^{\beta}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))$ (hence $\nu_{i}^{(1)}(\gamma_{\text{+}})$ is the normalizing constant for $\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))$). Further, the unnormalized conditional $\widetilde{\pi}_{i}(\cdot|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))$ at the next GS iteration differs from $\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))$ at no more than two points on its domain {-$\{\texttt{-}1\mathnormal{:}M\}$. Fig. 2 provides an illustration.
Figure 2: Difference between unnormalized conditionals in successive TGS iterations. If $\gamma^{\prime}_{\text{+}}$ differs from $\gamma_{\text{+}}$ only at the $n$-th coordinate, i.e., $\gamma_{\text{+}}(\ell_{n})=j>0$ and $\gamma^{\prime}_{\text{+}}(\ell_{n})=j^{\prime}>0$, with $j\neq j^{\prime}$, then for $i\in\{\bar{n}\}$, $\widetilde{\pi}_{i}(\cdot|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))$ is the same as $\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))$ on {-$\{\texttt{-}1\mathnormal{:}M\}$ except at $j$ (gray) and $j^{\prime}$ (red).
@@
Suppose that, $\gamma^{\prime}_{\text{+}}$ and $\gamma_{\text{+}}$ differ only at the $n$-th coordinate, i.e., $\gamma^{\prime}_{\text{+}}(\ell_{\bar{n}})=\gamma_{\text{+}}(\ell_{\bar{n}})$ and $\gamma^{\prime}_{\text{+}}(\ell_{n})\neq\gamma_{\text{+}}(\ell_{n})$. Then, for $i\in\{1\emph{:}P\}$, $j\in\{\texttt{-}1\emph{:}M\}$, and $\beta>0$,
@@
Proof.
See Appendix (Subsection VI-A). ∎
@@
Given successive TGS iterates $\gamma_{\text{+}}$ and $\gamma^{\prime}_{\text{+}}$ that differ at the $n$-th coordinate, Proposition 2 (with $\beta=1$) states that their $n$-th conditionals are the same, i.e., $\pi_{n}(\cdot|\gamma_{\text{+}}(\ell_{\bar{n}}))=\pi_{n}(\cdot|\gamma^{\prime}_{\text{+}}(\ell_{\bar{n}}))$. Moreover, for $i\neq n$, the unnormalized $i$-th conditional $\widetilde{\pi}_{i}(\cdot|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))$ is the same as $\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}(\ell_{\bar{i}}))$, except at:
  1. (i)
    $j\!=\gamma_{\text{+}}(\ell_{n})>0$, where $\eta_{i}(j)$ is added to $\widetilde{\pi}_{i}(j|\gamma_{\text{+}}(\ell_{\bar{i}}))$ so that $\widetilde{\pi}_{i}(j|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))=\eta_{i}(j)$ (because this $j$ is now freed up); and
  2. (ii)
    $j\!=\gamma^{\prime}_{\text{+}}(\ell_{n})>0$, where $-\eta_{i}(j)$ is added to $\widetilde{\pi}_{i}(j|\gamma_{\text{+}}(\ell_{\bar{i}}))$ so that $\widetilde{\pi}_{i}(j|\gamma^{\prime}_{\text{+}}(\ell_{\bar{i}}))=0$ (because this $j$ is now taken).
@@
Suppose that the successive TGS iterates $\gamma_{\text{+}}$ and $\gamma^{\prime}_{\text{+}}$ differ at the $n$-th coordinate. Then, for $\beta>0$,
@@
$\hskip 14.22636pt+\eta_{i}^{\beta}(\gamma_{\text{+}}(\ell_{n}))1_{\{1:M\}}(\gamma_{\text{+}}(\ell_{n}))$
$\hskip 14.22636pt-\eta_{i}^{\beta}(\gamma^{\prime}_{\text{+}}(\ell_{n}))1_{\{1:M\}}(\gamma^{\prime}_{\text{+}}(\ell_{n})).$
Input : $\gamma^{(0)}_{\text{+}}$, $T$, $\alpha$, $\beta$, $[\eta_{i}(j)]^{P}_{i\text{=}1}$, $\rho(\cdot|\gamma^{(0)}_{\text{+}})$, $[\widetilde{\pi}_{i}(\cdot|\gamma^{(0)}_{\text{+}}\!(\ell_{\bar{i}}))]^{P}_{i\text{=}1}$, $[\nu^{(1)}_{i}\!(\gamma^{(0)}_{\text{+}})]^{P}_{i\text{=}1}$, $[\nu^{(\beta)}_{i}\!(\gamma^{(0)}_{\text{+}})]^{P}_{i\text{=}1}$
Output : $[\gamma^{(t)}_{\text{+}}]^{T}_{t\text{=}1}$
1   $M\leftarrow\textsf{\small size}([\eta_{i}(j)]^{P}_{i\text{=}1},\textsf{\footnotesize col})\texttt{-}2$;
2 foreach $t=1:T$ do
3       $n\sim\textsf{\small Categorical}([1\mathnormal{:}P],\rho(\cdot|\gamma^{(t\text{-}1)}_{\text{+}}))$;
4       $\phi_{n}(\cdot|\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{\bar{n}}))\leftarrow\frac{\alpha\widetilde{\pi}_{n}(\cdot|\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{\bar{n}}))}{\nu^{(1)}_{n}\!(\gamma^{(t\text{-}1)}_{\text{+}})}+\frac{(1\text{-}\alpha)\widetilde{\pi}^{\beta}_{n}\!(\cdot|\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{\bar{n}}))}{\nu^{(\beta)}_{n}\!(\gamma^{(t\text{-}1)}_{\text{+}})}$;
5       $\gamma^{(t)}_{\text{+}}\!(\ell_{n})\sim\textsf{\small Categorical}([\texttt{-}1\mathnormal{:}M],\phi_{n}(\cdot|\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{\bar{n}})))$;
6       $\gamma^{(t)}_{\text{+}}\leftarrow[\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{1:n\text{-}1}),\gamma^{(t)}_{\text{+}}\!(\ell_{n}),\gamma^{(t\text{-}1)}_{\text{+}}\!(\ell_{n\text{+}1:P})]$;
7       foreach $i=1:P$ do
8             Compute $\widetilde{\pi}_{i}(\cdot|\gamma^{(t)}_{\text{+}}\!(\ell_{\bar{i}})),\nu^{(1)}_{i}\!(\gamma^{(t)}_{\text{+}}),\nu^{(\beta)}_{i}\!(\gamma^{(t)}_{\text{+}})$ +++++++++++++++++++++++++++ via Corollary 1;
9             $\pi_{i}(\gamma_{\text{+}}^{(t)}\!(\ell_{i})|\gamma^{(t)}_{\text{+}}\!(\ell_{\bar{i}}))\leftarrow\frac{\widetilde{\pi}_{i}(\gamma_{\text{+}}^{(t)}\!(\ell_{i})|\gamma^{(t)}_{\text{+}}\!(\ell_{\bar{i}}))}{\nu^{(1)}_{i}\!(\gamma^{(t)}_{\text{+}})}$;
10             $\phi_{i}(\gamma_{\text{+}}^{(t)}\!(\ell_{i})|\gamma^{(t)}_{\text{+}}\!(\ell_{\bar{i}}))\leftarrow\alpha\pi_{i}(\gamma_{\text{+}}^{(t)}\!(\ell_{i})|\gamma^{(t)}_{\text{+}}\!(\ell_{\bar{i}}))$ $\phantom{++++++++++++}+\frac{(1\text{-}\alpha)\widetilde{\pi}^{\beta}_{i}\!(\gamma_{\text{+}}^{(t)}\!(\ell_{i})|\gamma^{(t)}_{\text{+}}\!(\ell_{\bar{i}}))}{\nu^{(\beta)}_{i}\!(\gamma^{(t)}_{\text{+}})}$;
11             $\widetilde{\rho}(i|\gamma^{(t)}_{\text{+}})\leftarrow\frac{\phi_{i}(\gamma_{\text{+}}^{(t)}\!(\ell_{i})|\gamma^{(t)}_{\text{+}}\!(\ell_{\bar{i}}))}{\pi_{i}(\gamma_{\text{+}}^{(t)}\!(\ell_{i})|\gamma^{(t)}_{\text{+}}\!(\ell_{\bar{i}}))}$;
12
13      $\rho(\cdot|\gamma^{(t)}_{\text{+}})\leftarrow\frac{\widetilde{\rho}(\cdot|\gamma^{(t)}_{\text{+}})}{\langle\widetilde{\rho}(\cdot|\gamma^{(t)}_{\text{+}}),1\rangle}$;
14
Algorithm 2 $TGS^{+}$
@@
The above result means propagating each tempered conditional and its normalizing constant to the next iterate of the Markov chain can be performed with a constant time complexity. Consequently, $\mathcal{O}(T(P+M))$ complexity TGS can be developed for selecting significant GLMB components. The pseudocode in Algorithm 2, herein referred to as $TGS^{+}$, outlines the steps for generating iterates $\gamma_{\text{+}}^{(1)},\cdots,\gamma_{\text{+}}^{(T)}$, from initial chain state $\gamma_{\text{+}}^{(0)}$, and initial coordinate distribution $\rho(\cdot|\gamma_{\text{+}}^{(0)})$. Note that:
@@
  • Sampling $n$ from the categorical distribution $\rho(\cdot|\gamma_{\text{+}}^{(t\text{-}1)})$, defined on $P$ categories, incurs $\mathcal{O}(P)$ complexity (line 2);
@@
  • Computing the proposal $\phi_{n}(\cdot|\gamma_{\text{+}}^{(t\text{-}1)}(\ell_{\bar{n}}))$ incurs $\mathcal{O}(M)$ complexity (line 2) since $\widetilde{\pi}_{i}(\cdot|\gamma_{\text{+}}^{(t\text{-}1)}(\ell_{\bar{i}}))$, $\nu_{i}^{(1)}(\gamma_{\text{+}}^{(t\text{-}1)})$, and $\nu_{i}^{(\beta)}(\gamma_{\text{+}}^{(t\text{-}1)})$ were generated as by-products of computing $\rho(\cdot|\gamma_{\text{+}}^{(t\text{-}1)})$ in the previous iteration;
@@
  • Sampling $\gamma_{\text{+}}^{(t)}(\ell_{n})$ from $\phi_{n}(\cdot|\gamma_{\text{+}}^{(t\text{-}1)}(\ell_{\bar{n}}))$ incurs $\mathcal{O}(M)$ complexity (line 2);
@@
  • Computing $\widetilde{\rho}(\cdot|\gamma_{\text{+}}^{(t)})$ for the $(t+1)$-th iteration is only needed when $\gamma_{\text{+}}^{(t)}(\ell_{n})\neq\gamma_{\text{+}}^{(t\texttt{-}1)}(\ell_{n})$, and requires $\mathcal{O}(P)$ complexity (lines 2-2), because for each $i\in\{1\mathnormal{:}P\}$, evaluating (19) from the available $\pi_{i}(\gamma_{\text{+}}^{(t)}(\ell_{i})|\gamma_{\text{+}}^{(t)}(\ell_{\bar{i}}))$ and $\phi_{i}(\gamma_{\text{+}}^{(t)}(\ell_{i})|\gamma_{\text{+}}^{(t)}(\ell_{\bar{i}}))$ only requires a constant time complexity (Corollary 1); and